A helium balloon can just lift a load of 800 N. The skin of the balloon has a mass of 1.10 kg. What is the volume of the balloon?(the density of air is 1.29 kg/m³ and the density of helium is 0.179 kg/m³)
Answer #1

here

F(archimedes) - w(load) - w(balloon) - w(helium) = 0

p(air) * g * V - w(load) - w(balloon) - p(helium) * g * V = 0

1.29 * 9.8 * V - 800 - 1.1 * 9.8 - 0.179 * 9.8 * V = 0

V = 74.46 m^3